# Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface

### v2

@@ II-B Problem Formulation
where ${\bf w}\triangleq[{\bf w}_{1}^{H},\ldots,{\bf w}_{G}^{H}]^{H}$, and $\gamma_{ik}$ is the SINR target for user $k$ in group $i$.
Problem ${\cal P}_{o}$ is a non-convex quadratically constrained quadratic programming (QCQP) problem. Note that for the conventional downlink transmission without RIS, the multicast beamforming problem is already known to be NP-hard in its general form. The joint optimization problem ${\cal P}_{o}$ is even more challenging with the additional RIS passive beamformer ${\bf e}$ affecting $\textrm{SINR}_{ik}$ to be optimized under the non-convex constraints in (3c) imposed. Furthermore, since the number of RIS elements $M$ is expected to be large to show benefit, ${\cal P}_{o}$ is typically a large-scale problem. Thus, our goal is to develop an effective computational method to provide a solution with high performance but low computational complexity.
@@ III RIS-assisted Multicast Beamforming Design
We consider the alternating optimization (AO) technique to solve ${\cal P}_{o}$, i.e., solving ${\cal P}_{o}$ with respect to (w.r.t.) the BS multicast beamformer ${\bf w}$ and RIS passive beamformer ${\bf e}$ alternatingly. However, directly applying AO to ${\cal P}_{o}$ is not an effective method. This is because the objective function in ${\cal P}_{o}$ is only a function of ${\bf w}$. With given ${\bf w}$, ${\cal P}_{o}$ is reduced to a feasibility problem w.r.t. ${\bf e}$, for which the value of ${\bf e}$ from the previous AO iteration is already a feasible solution. Thus, the algorithm could get stuck at the value of $({\bf w},{\bf e})$ after the first round of updates, and the quality of this solution merely depends on the quality of the initial point.
In our approach, we first analyze the structure of ${\cal P}_{o}$ w.r.t. ${\bf w}$ and ${\bf e}$ and then utilize it to devise an effective algorithm to solve this QoS problem.
III-A QoS Problem Reformulation
@@
We can replace (3b) with (4) and equivalently rewrite ${\cal P}_{o}$ as
@@
s.t. $\ \ \max_{{\bf e}:|e_{m}|^{2}=1,m\in{\cal M}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1.$ (6b)
Note that for showing the equivalence, it is straightforward to show ${\cal P}_{1}$ and ${\cal P}_{2}$ have the same feasible set.
Comparing ${\cal P}_{2}$ with ${\cal P}_{o}$, we see that the original set of SINR constraints in (3b) is replaced by the constraint in (6b). The left-hand side of (6b) is a max-min weighted SINRs optimization problem w.r.t. ${\bf e}$, and the unit-modulus constraints on ${\bf e}$ in (3c) are now the constraints for this max-min optimization problem.
III-B Alternating Multicast Beamforming Approach
Once transforming ${\cal P}_{o}$ into ${\cal P}_{2}$, we notice that ${\cal P}_{2}$ inherently contains two optimization subproblems w.r.t. ${\bf w}$ and ${\bf e}$, respectively. As we will show below, these two subproblems are essentially two types of multicast beamforming problems: 1) BS multicast beamforming QoS problem w.r.t. ${\bf w}$; 2) RIS passive multicast beamforming MMF problem w.r.t. ${\bf e}$. Following this, we naturally break down ${\cal P}_{2}$ into two subproblems to solve alternatingly. Our overall proposed alternating multicast beamforming (AMBF) approach includes two steps: the alternating optimization step and a final processing step. The final processing step is to ensure a feasible solution to ${\cal P}_{o}$. Below, we will first describe the two subproblems and the final processing step, leaving the fast algorithm design for solving the subproblems to Sections IV-A and IV-B.
@@ III-B1 BS Multicast Beamforming QoS Problem for ${\bf w}$
For given RIS passive beamforming vector ${\bf e}$, ${\cal P}_{o}$ is reduced to the following BS multicast beamforming problem w.r.t. ${\bf w}$:
@@
s.t. $\ \ \textrm{SINR}_{ik}\geq\gamma_{ik},\ k\in{\cal K}_{i},~{}i\in{\cal G}.$ (7b)
Problem ${\cal P}_{\bf e}$ is a typical downlink multi-group multicast beamforming QoS problem. The effective channel between the BS and user $k$ in group $i$ consists of the direct path and the RIS-path given by ${\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}$, as shown in $\textrm{SINR}_{ik}$ in (2).
@@ III-B2 RIS Passive Multicast Beamforming MMF Problem for ${\bf e}$
Optimizing RIS passive beamformer ${\bf e}$ with given ${\bf w}$ is the main challenge in ${\cal P}_{o}$. As discussed earlier, directly applying AO to the original problem ${\cal P}_{o}$ is not effective. However, once we convert ${\cal P}_{o}$ into ${\cal P}_{2}$, for given ${\bf w}$, we naturally have the following optimization problem w.r.t. ${\bf e}$, which is to maximize the minimum ratio of the received SINR over its target, or the weighted SINR, i.e., ${\textrm{SINR}_{ik}}/{\gamma_{ik}}$:
@@
s.t. $\ \ |e_{m}|^{2}=1,\ m\in{\cal M}.$ (8b)
@@
Note that ${\cal S}_{{\bf w}}$ can be viewed as a weighted MMF single-group multicast beamforming problem, with ${\bf e}$ being the beamformer to all the users in the system. However, this problem has three key differences from the conventional MMF multicast problem: i) This single-group multicast beamforming problem contains self-interference, as ${\bf e}$ appears in the denominator of $\textrm{SINR}_{ik}$ in (2). ii) From the point of view of optimizing ${\bf e}$, $\textrm{SINR}_{ik}$ in (2) can in fact be interpreted as the signal-to-leakage-and-noise ratio (SLNR) for user $k$ in group $i$. Thus, the MMF problem ${\cal S}_{{\bf w}}$ w.r.t. ${\bf e}$ is based on an SLNR metric. iii) Different from the conventional transmit power constraint, the constraints in (8b) can be viewed as a non-convex per-element power constraint, where each element is required to consume exactly one unit of power.
Due to the difference discussed above, ${\cal S}_{\bf w}$ is a more challenging problem to solve than the conventional MMF problem, which is a non-convex QCQP problem that is NP-hard. To tackle this problem, we note that the constraints in (8b) are equivalent to
$|e_{m}|^{2}\leq 1,m\in{\cal M};\quad\quad{\bf e}^{T}{\bf e}=M.$ (9)
@@
To make the problem more tractable, we relax ${\cal S}_{\bf w}^{\text{eq}}$ by transferring the constraint in (10c) into the objective function in (10a) as a penalty term with a penalty weight $\zeta>0$. The relaxed problem is given by
$\widetilde{{\cal S}}_{{\bf w}}:\ \max_{{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\zeta\frac{{\bf e}^{H}{\bf e}}{M}$ (11a)
@@
The penalty term in (11a) is to ensure that ${\bf e}^{T}{\bf e}$ is as close to $M$ as possible in solving $\widetilde{{\cal S}}_{{\bf w}}$. Note that we choose the penalty term in this form such that the two terms in the objective function are both normalized. In particular, for the weighted SINR term, using $({\bf w},{\bf e})$ from ${\cal P}_{\bf e}$, we have $\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}=1$ based on (7b). Thus, we expect this term to be close to $1$, especially as the AO iteration between ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{\bf w}$ converges. The RIS passive beamformer norm term satisfies $\frac{{\bf e}^{H}{\bf e}}{M}\leq 1$ based on (10b), and it is also expected to be close to $1$ as the AO iteration converges. In this way, the two terms in (11a) have comparable values that are both around $1$, which are not affected by the actual values of SINR or $\|{\bf e}\|^{2}$ for a specific setup. This simplifies the tuning of penalty parameter $\zeta$, which can remain the same for different system configurations.
Although $\widetilde{{\cal S}}_{{\bf w}}$ is still a non-convex QCQP problem, compared with ${\cal S}_{{\bf w}}$, it is more amenable to efficient algorithm design to compute a solution, which will be discussed in Section IV-B.
Thus, following the above, the first step in our AMBF approach is to solve the two multicast beamforming problems alternatingly, i.e., ${\cal P}_{\bf e}$ for the BS beamformers $\{{\bf w}_{i}\}$ and $\widetilde{{\cal S}}_{{\bf w}}$ for the RIS beamformer ${\bf e}$.
@@ III-B3 Final Processing
Let $({\bf w}^{\star},{\bf e}^{\star})$ denote the solution after the above iterative procedure. Since $\widetilde{{\cal S}}_{{\bf w}}$ is a relaxed problem of ${\cal S}_{{\bf w}}$, $({\bf w}^{\star},{\bf e}^{\star})$ may not be feasible to ${\cal P}_{o}$. Thus, based on $({\bf w}^{\star},{\bf e}^{\star})$, we have the final processing step to obtain the feasible solution $({\bf w}^{\text{final}},{\bf e}^{\text{final}})$ as follows:
  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which represents taking the phase component of each element in ${\bf e}^{\star}$.
@@
IV Fast Algorithms for the AMBF Approach
In this section, we propose our fast algorithms to solve subproblems ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{{\bf w}}$.
IV-A Fast Algorithm for BS Multicast Beamforming ${\cal P}_{\bf e}$
As mentioned earlier, ${\cal P}_{\bf e}$ is a typical downlink multicast beamforming QoS problem. Define $\tilde{{\bf h}}_{ik}\triangleq{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}$ as the $N\times 1$ equivalent channel from BS to user $k$ in group $i$ that includes both the direct path and the RIS path. The received SINR at user $k$ in group $i$ in (2) can be rewritten as
$\textrm{SINR}_{ik}=\frac{|{\bf w}_{i}^{H}\tilde{{\bf h}}_{ik}|^{2}}{\sum_{j\neq i}^{G}|{\bf w}_{j}^{H}\tilde{{\bf h}}_{ik}|^{2}+\sigma^{2}},\ k\in{\cal K}_{i},i\in{\cal G}.$ (12)
Although the QoS problem ${\cal P}_{\bf e}$ is known to be a challenging non-convex QCQP problem, the structure of the optimal solution is obtained in [11], which can be utilized to compute the beamforming solution ${\bf w}$ with high computational efficiency, especially when the number of BS antennas is large $M\gg 1$. We directly employ this optimal structure to compute ${\bf w}$. Specifically, the optimal ${\bf w}_{i}$ is a weighted MMSE beamforming given by [11]
@@ IV-A Fast Algorithm for BS Multicast Beamforming ${\cal P}_{\bf e}$
where $\widetilde{{\bf H}}_{i}\triangleq[\tilde{{\bf h}}_{i1},\ldots,\tilde{{\bf h}}_{iK}]$ is the equivalent channel matrix for group $i$, ${\bf a}_{i}$ is the $K\times 1$ optimal weight vector for group $i$ containing the (complex) weight for each user channel in the group, and ${\bf R}(\boldsymbol{\lambda})\triangleq{\bf{I}}+\sum_{i=1}^{G}\sum_{k=1}^{K}\lambda_{ik}\gamma_{ik}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}$ is the noise-plus-weighted-channel covariance matrix with $\lambda_{ik}$ being the optimal Lagrange multipliers associated with the SINR constraint in (7b) and $\boldsymbol{\lambda}$ being the vector containing all $\lambda_{ik}^{o}$’s.
The solution ${\bf w}_{i}$ in (13) is given in a semi-closed form with $\boldsymbol{\lambda}$ and $\{{\bf a}_{i}\}$ to be numerically determined. Although the optimal value for $\boldsymbol{\lambda}$ is difficult to obtain, an algorithm to compute (suboptimal) values of $\boldsymbol{\lambda}$ efficiently using the fixed-point method is proposed in [11]. With the optimal structure of ${\bf w}_{i}$ in (13) and $\boldsymbol{\lambda}$, ${\cal P}_{\bf e}$ can be transformed into a weight optimization problem w.r.t. weight vectors $\{{\bf a}_{i}\}$, which is a much smaller problem with $\sum_{i=1}^{G}K_{i}$ variables compared to the original problem ${\cal P}_{\bf e}$ with $GN$ variables for $N\gg K_{i}$. To solve this non-convex weight optimization problem for $\{{\bf a}_{i}\}$, we can adopt the SCA approach discussed in [11]. Further improvement in the computational efficiency can be considered by adopting the fast ADMM algorithm proposed recently in [13] based on the optimal structure, which provides each SCA update using closed-form expressions with ultra-low complexity. We omit the details and refer the readers to [11, 13] for these algorithms.
IV-B Fast Algorithm for RIS Passive Multicast Beamforming $\widetilde{{\cal S}}_{\bf w}$
Problem $\widetilde{{\cal S}}_{{\bf w}}$ is expected to be large-size, since the number of RIS elements is large ($M\gg 1$). Existing methods in the literature for solving ${\bf e}$ often adopt SDR or SCA [17, 20], but suffer from high computational complexity. It is essential to develop an effective low-complexity algorithm to find a solution to $\widetilde{{\cal S}}_{{\bf w}}$, especially when this subproblem needs to be solved in each AO iteration.
From Remark 1, we note that $\widetilde{{\cal S}}_{{\bf w}}$ is a variant of the MMF problem. PSA is a fast first-order algorithm recently proposed in[12] to solve the multi-group multicast beamforming MMF problem in the conventional system. Although the form of the objective function and the constraints in $\widetilde{{\cal S}}_{{\bf w}}$ are different from those of the conventional MMF problem considered in [12], the method can still be adapted to solve $\widetilde{{\cal S}}_{{\bf w}}$. Thus, we propose to apply PSA to $\widetilde{{\cal S}}_{{\bf w}}$ to compute a near-stationary solution of $\widetilde{{\cal S}}_{{\bf w}}$ efficiently.

### v3

@@ II-B Problem Formulation
where ${\bf w}\triangleq[{\bf w}_{1}^{H},\ldots,{\bf w}_{G}^{H}]^{H}$, and $\gamma_{ik}$ is the SINR target for user $k$ in group $i$. Note that the SINR constraint in (2b) is equivalent to the per-user minimum rate constraint; it essentially defines a QoS target for each user. When $\gamma_{ik}=\tilde{\gamma}_{i}$, $\forall k$, this target becomes the group rate target, and constraint (2b) reduce to ensure that the minimum group rate target $\tilde{\gamma}_{i}$ is met for group $i$.
Problem ${\cal P}_{o}$ is a non-convex quadratically constrained quadratic programming (QCQP) problem. For conventional downlink transmission without RIS, the multicast beamforming problem w.r.t. ${\bf w}$ is known to be NP-hard. Therefore, the joint optimization problem ${\cal P}_{o}$ is even more challenging to solve, since it includes the additional RIS passive beamformer ${\bf e}$ affecting $\textrm{SINR}_{ik}$ to be optimized under the non-convex unit-modulus constraint (2c). Furthermore, the number of RIS elements $M$ is expected to be large to realize RIS benefits. Thus, ${\cal P}_{o}$ is typically a large-scale problem to solve. Our goal is to develop a fast and scalable algorithmic solution that is of high performance for RIS-assisted multicast beamforming.
@@ III RIS-assisted Multicast Beamforming Design
A common approach to solve ${\cal P}_{o}$ is to consider using alternating optimization (AO) to compute the BS multicast beamformer ${\bf w}$ and RIS passive beamformer ${\bf e}$ alternatingly. However, directly applying AO to ${\cal P}_{o}$ is not an effective method. This is because the objective is only a function of ${\bf w}$. Given ${\bf w}$, ${\cal P}_{o}$ is reduced to a feasibility problem w.r.t. ${\bf e}$, which cannot ensure the quality of solution. For example, the value of ${\bf e}$ from the previous AO iteration is already a feasible solution. The algorithm could stuck at the value of $({\bf w},{\bf e})$ after the first round of updates, and the quality of this solution merely depends on the quality of the initial point. Thus, instead of directly applying AO, we first analyze the structure of ${\cal P}_{o}$ w.r.t. ${\bf w}$ and ${\bf e}$ and then utilize it to devise an effective algorithm to solve ${\cal P}_{o}$.
III-A Reformulation
@@
Replacing (2b) with (3), we can equivalently rewrite ${\cal P}_{o}$ as
@@
s.t. $\ \ \max_{\begin{subarray}{c}{\bf e}:|e_{m}|^{2}=1\\ m\in{\cal M}\end{subarray}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1.$ (5b)
To see the equivalence, it is straightforward to show ${\cal P}_{1}$ and ${\cal P}_{2}$ have the same feasible set.
From ${\cal P}_{o}$ to ${\cal P}_{2}$, the original set of SINR constraints in (2b) is now replaced by the constraint (5b). Note that the left hand side of this new constraint is a max-min weighted SINRs optimization problem w.r.t. ${\bf e}$, and the unit-modulus constraints on ${\bf e}$ in (2c) are now the constraints for this max-min optimization problem. Thus, the joint optimization problem ${\cal P}_{o}$ is transformed into ${\cal P}_{2}$, which inherently contains two separate optimization subproblems w.r.t. ${\bf w}$ and ${\bf e}$, respectively. Following this, we can naturally break down ${\cal P}_{2}$ into two subproblems to solve alternatingly.
III-B Proposed Alternating Multicast Beamforming Approach
Examining the two subproblems in ${\cal P}_{2}$, we will show below that they are essentially two types of multicast beamforming problems: 1) BS multicast beamforming QoS problem w.r.t. ${\bf w}$; 2) RIS passive multicast beamforming MMF problem w.r.t. ${\bf e}$. We solve these two subproblems alternatingly. Our proposed alternating multicast beamforming (AMBF) approach includes two steps: 1) the alternating optimization step, and 2) a final processing step. The final processing step is to ensure a feasible solution to ${\cal P}_{o}$. Below, we will focus on describing the two subproblems and the final processing step, leaving the algorithm design for solving the subproblems to Sections IV-A and IV-B.
@@ III-B1 BS Multicast Beamforming QoS Problem for ${\bf w}$
Given RIS passive beamforming vector ${\bf e}$, ${\cal P}_{o}$ is reduced to the following BS multicast beamforming problem w.r.t. ${\bf w}$:
@@
s.t. $\ \ \textrm{SINR}_{ik}\geq\gamma_{ik},\ k\in{\cal K}_{i},~{}i\in{\cal G},$ (6b)
which is a conventional downlink multi-group multicast beamforming QoS problem. The effective channel between the BS and user $k$ in group $i$ consists of the direct path and the RIS-path given by ${\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}$, as shown in $\textrm{SINR}_{ik}$ in (1).
@@ III-B2 RIS Passive Multicast Beamforming MMF Problem for ${\bf e}$
Given ${\bf w}$, RIS passive beamformer ${\bf e}$ optimization is the main challenge in ${\cal P}_{o}$. However, once we convert ${\cal P}_{o}$ into ${\cal P}_{2}$, we see that optimizing ${\bf e}$ is essentially to maximize the minimum ratio of the received SINR at each user over its target, i.e., the weighted SINR, as the following max-min problem:
@@
s.t. $\ \ |e_{m}|^{2}=1,\ m\in{\cal M}$ (7b)
where $\textrm{SINR}_{ik}$ in (1) can be re-expressed as
$\textrm{SINR}_{ik}=\frac{|{\bf e}^{H}{\bf G}_{ik}^{H}{\bf w}_{i}+{\bf h}^{\text{d}}_{ik}\!{}^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}|{\bf e}^{H}{\bf G}_{ik}^{H}{\bf w}_{j}+{\bf h}^{\text{d}}_{ik}\!{}^{H}{\bf w}_{j}|^{2}+\sigma^{2}}.$ (8)
@@
Note that ${\cal S}_{{\bf w}}$ can be viewed as a variant of single-group multicast beamforming weighted MMF problem, with ${\bf e}$ being the multicast beamformer to all the users. This can be more clearly seen in the RIS-path only case, given by
$\textrm{SINR}_{ik}=\frac{|{\bf e}^{H}{\bf G}_{ik}^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}|{\bf e}^{H}{\bf G}_{ik}^{H}{\bf w}_{j}|^{2}+\sigma^{2}}.$ (9)
But it has a few differences from the conventional MMF multicast beamforming: i) It contains self-interference, since ${\bf e}$ appears in the denominator of $\textrm{SINR}_{ik}$ in (9). ii) From the perspective of RIS beamformer ${\bf e}$, $\textrm{SINR}_{ik}$ in (9) can be interpreted as the signal-to-leakage-and-noise ratio (SLNR) for user $k$ in group $i$. The power leakage is through the effective channels formed by other BS beamformers ${\bf w}_{j}$’s to the same user, ${\bf G}_{ik}^{H}{\bf w}_{j}$, $\forall j\neq i$. Thus, the MMF problem ${\cal S}_{{\bf w}}$ w.r.t. ${\bf e}$ is based on an SLNR metric. For the general case with a direct path in (8), the combined beamed signal strength from the RIS and the direct path need to be considered. iii) Different from the conventional total transmit power constraint, the constraints in (7b) can be viewed as a non-convex per-element power constraint, where each element is required to consume exactly one unit of power.
Due to the difference discussed above, ${\cal S}_{\bf w}$ is a more challenging problem to solve than the conventional MMF problem, which is a non-convex QCQP problem. To tackle this problem, we note that the constraints in (7b) are equivalent to
$|e_{m}|^{2}\leq 1,m\in{\cal M}\quad\text{and}\quad{\bf e}^{T}{\bf e}=M.$ (10)
@@
To make the problem more tractable, we relax ${\cal S}_{\bf w}^{\text{eq}}$ by transferring the constraint in (11c) into the objective function in (11a) as a penalty term with a penalty weight $\rho>0$. The relaxed problem is given by
$\widetilde{{\cal S}}_{{\bf w}}:\ \max_{{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\rho\frac{{\bf e}^{H}{\bf e}}{M}$ (12a)
@@
The penalty term in (12a) is to ensure that ${\bf e}^{T}{\bf e}$ is as close to $M$ as possible in solving $\widetilde{{\cal S}}_{{\bf w}}$. Note that we choose the penalty term in this particular form such that the two terms in the objective function are both normalized. In particular, for first term, we note that ${\bf w}$ is the solution to ${\cal P}_{\bf e}$, which yields $\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}=1$ in (6b). Thus, we expect this term to be close to $1$, especially when the iterative procedure between ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{\bf w}$ converges. The second term for the normalized RIS passive beamformer satisfies $\frac{{\bf e}^{H}{\bf e}}{M}\leq 1$ based on (12b), and it is also expected to be close to $1$ when the AO procedure converges. Therefore, the two terms in (12a) have comparable values that are both close to $1$. Being normalized, they do not depend on the actual values of SINR or $\|{\bf e}\|^{2}$ for a specific system configuration. This simplifies the tuning of penalty parameter $\rho$, which can remain the same for different system configurations.
Compared with ${\cal S}_{{\bf w}}$, $\widetilde{{\cal S}}_{{\bf w}}$ is more amenable to efficient algorithm design to compute a solution, which will be discussed in Section IV-B.
Combining Sections III-B1 and III-B2, the first step in our AMBF approach is to solve the two multicast beamforming problems alternatingly, i.e., ${\cal P}_{\bf e}$ for the BS beamformers $\{{\bf w}_{i}\}$ and $\widetilde{{\cal S}}_{{\bf w}}$ for the RIS beamformer ${\bf e}$.
@@ III-B3 Final Processing
Let $({\bf w}^{\star},{\bf e}^{\star})$ denote the solution from the above AO procedure. Since $\widetilde{{\cal S}}_{{\bf w}}$ is a relaxed problem of ${\cal S}_{{\bf w}}$, $({\bf w}^{\star},{\bf e}^{\star})$ may not be feasible to ${\cal P}_{o}$. Thus, based on $({\bf w}^{\star},{\bf e}^{\star})$, we have the final processing step to obtain the feasible solution $({\bf w}^{\text{final}},{\bf e}^{\text{final}})$ as follows:
  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which is taking the phase component of each element in ${\bf e}^{\star}$.
@@
IV Fast Algorithms for alternating multicast beamforming
We now propose our fast algorithm for each of the multicast subproblems ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{{\bf w}}$ under the proposed AMBF approach.
IV-A BS Multicast Beamforming QoS Problem ${\cal P}_{\bf e}$
As mentioned earlier, ${\cal P}_{\bf e}$ is a conventional downlink multicast beamforming QoS problem. Define $\tilde{{\bf h}}_{ik}\triangleq({\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e})$ as the $N\times 1$ equivalent channel from BS to user $k$ in group $i$ that includes both the direct path and the RIS path. The received SINR in (1) at user $k$ in group $i$ can be rewritten as
$\textrm{SINR}_{ik}=\frac{|\tilde{{\bf h}}_{ik}^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}^{G}|\tilde{{\bf h}}_{ik}^{H}{\bf w}_{j}|^{2}+\sigma^{2}},\ k\in{\cal K}_{i},i\in{\cal G}.$ (13)
Despite that ${\cal P}_{\bf e}$ is an NP-hard non-convex QCQP problem, the structure of the optimal solution is obtained in [20], which can be utilized to compute the beamforming solution ${\bf w}$ with high computational efficiency, especially when the number of BS antennas is large $N\gg 1$. We directly employ this optimal structure to compute ${\bf w}$. Specifically, the optimal ${\bf w}_{i}$ is a weighted MMSE beamforming given by [20]
@@ IV-A BS Multicast Beamforming QoS Problem ${\cal P}_{\bf e}$
where $\widetilde{{\bf H}}_{i}\triangleq[\tilde{{\bf h}}_{i1},\ldots,\tilde{{\bf h}}_{iK}]$ is the equivalent channel matrix for users in group $i$, ${\bf a}_{i}$ is the $K\times 1$ optimal weight vector for group $i$ containing the (complex) weight for each user channel in the group, and ${\bf R}(\boldsymbol{\lambda})\triangleq{\bf{I}}+\sum_{i=1}^{G}\sum_{k=1}^{K}\lambda_{ik}\gamma_{ik}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}$ is the noise-plus-channel covariance matrix with $\lambda_{ik}$ being the optimal Lagrange multipliers associated with the SINR constraint in (6b) and $\boldsymbol{\lambda}$ being the vector containing all $\lambda_{ik}^{o}$’s.
The solution ${\bf w}_{i}$ in (14) is given in a semi-closed form with $\boldsymbol{\lambda}$ and $\{{\bf a}_{i}\}$ to be numerically determined. We can use the algorithm proposed in [20] to compute (suboptimal) values of $\boldsymbol{\lambda}$ efficiently using the fixed-point method. Given $\boldsymbol{\lambda}$ and ${\bf w}_{i}$ in (14), ${\cal P}_{\bf e}$ can be transformed into a weight optimization problem w.r.t. weight vectors $\{{\bf a}_{i}\}$, which is a much smaller problem with $\sum_{i=1}^{G}K_{i}$ variables compared to the original problem ${\cal P}_{\bf e}$ with $GN$ variables for $N\gg K_{i}$. To solve the non-convex weight optimization problem for $\{{\bf a}_{i}\}$, we can adopt the fast algorithm proposed in [22]. It is based on the SCA approach and uses an ADMM-based algorithm to provide each SCA update using closed-form expressions with ultra-low complexity. We omit the details and refer the readers to [20, 22] for these algorithms.
IV-B RIS Passive Multicast Beamforming MMF Problem $\widetilde{{\cal S}}_{\bf w}$
The max-min problem $\widetilde{{\cal S}}_{{\bf w}}$ is expected to be large-size, since the number of RIS elements $M$ is large. Existing methods for solving ${\bf e}$ often adopt certain convexification methods, such as SDR or SCA [27, 30]. However, they suffer from high computational complexity for large $M$. Given this subproblem needs to be solved in each AO iteration., it is essential to develop an effective low-complexity algorithm to find a solution to $\widetilde{{\cal S}}_{{\bf w}}$.
From Remark 1, we note that $\widetilde{{\cal S}}_{{\bf w}}$ is a variant of the MMF problem. We propose to apply the projected subgradient algorithm (PSA) [31] to $\widetilde{{\cal S}}_{{\bf w}}$. PSA is a first-order iterative algorithm. It has been adopted in[21] to solve the conventional multi-group multicast beamforming MMF problem efficiently. The structure of the objective function and constraints in $\widetilde{{\cal S}}_{\bf w}$ are different from the conventional MMF problem. Nonetheless, we can still employ the essential technique of this method to compute a near-stationary solution of $\widetilde{{\cal S}}_{{\bf w}}$ efficiently. Given our problem structure, applying PSA proves to be computationally inexpensive, which is the main advantage of this approach.